OpenStudy (anonymous):

how do I finish this problem? X²+2+2x=x-2+x²+1 -x -x --------------------------------- 2+2x=x-2+1 2+2x=x-1 How do I finish this problem?

OpenStudy (anonymous):

Give me a guess.

OpenStudy (anonymous):

Get x on one side. 2+ 2x - x = x - 1 - x subtracting by x on both sides will get u this

OpenStudy (anonymous):

So x + 2 = -1 x = -3

OpenStudy (anonymous):

Nevermind lol. He got it

OpenStudy (anonymous):

oh my God thank you so much @master50777

OpenStudy (anonymous):

anytime